cesar is now 8 times as old as his son in 4 years he will be 4 times as as his son. how old is he.
Answer by mamiya(56)

let c be Cesar's age and s his son's.

Cesar is now 8 times as old as his son, means c = 8s
in 4 years, he will be 4 times as his son, means c+4= 4(s+4)
c + 4= 4s + 16
8s + 4= 4s + 16
4s = 12

s=3, so c=8*s=8*3=24
so Cesar is 24 and his son is 3.